OpenStudy (anonymous):

help please!! 1 medal and fan 6. If m

Directrix (directrix):

In a triangle, the largest side is opposite the largest angle of the triangle. Look at the attached diagram. The smallest angle is L. What side of the triangle is opposite angle L? @knowel

OpenStudy (anonymous):

true

OpenStudy (anonymous):

so its c

OpenStudy (anonymous):

or b

Directrix (directrix):

Side. What side is opposite angle L. A side is a segment, not a vertex. The opposite side of an angle does not form the angle.

Directrix (directrix):

Look at this.

Directrix (directrix):

#8 The midsegment, also known as the segment joining two sides of a triangle has measure half that of the third side of the triangle. So, solve 5x +2 = (1/2)* (3x + 32) for x. I would begin by doubling both sides. 2*(5x + 2) = (3x + 32) @knowel Solve this for x. Then substitute the value of x you get into 5x + 2 which is the algebraic value of the midsegment.

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

22

Directrix (directrix):

#9 The first step in an indirect proof is to assume the negation of what you are trying to prove. That means to insert a "not" in this statement: The figure is a triangle. @knowel

OpenStudy (anonymous):

c because you said not right?

Directrix (directrix):

#10 Maria passed the English exam. The negation of this statement involves the insertion of a "not" into the statement. Note: A statement and its negation cannot be simultaneously true. @knowel

OpenStudy (anonymous):

10. c

Directrix (directrix):

@knowel Without the problem numbers on your posts, I don't know which problem you are doing. That's another reason it is a good idea to post exactly one question per thread. What did you get for #6?

OpenStudy (anonymous):

ok so 6.c ,

OpenStudy (anonymous):

7.b 8.a 9.c 10.c

Directrix (directrix):

And, here's another issue. You do not list the letters of the answers in the posts yet you answer with the letter of the option. So, that means a helper has to scroll back up the thread and try to figure out which option, in this case, "c" is. So, please make it easier on your helpers. Rather than cite the answers by letters which are not given with the options, cite the total answer OR include the letters of the options with the problem posts. @knowel

OpenStudy (anonymous):

ok

Directrix (directrix):

No clue on what answers these letters go with so I cannot check. 7.b 8.a 9.c 10.c I don't know.

OpenStudy (anonymous):

7.<VWZ congruent <YWZ 8.22 9.Assume the figure is not a triangle 10.Maria did not take the English exam.

Directrix (directrix):

10.Maria did not take the English exam. Not correct. Let's move on. ONE question per thread. Write the letters of the answer options OR when you give your answer, post the entire answer. @knowel